Echo Tech. announced today that it is changing is dividend policy in an effort to reinvest in itself. The firm just paid a $3.00 dividend, which is expected to decline by 10% next year and decline another 20% two years from today. The dividend is then expected to grow at a constant 5% rate thereafter. Find the value of the stock today if investors require a 15% rate of return. Round intermediate steps to four decimals.
Answer #1

D1=(3*(1-0.1)=2.7

D2=(2.7*(1-0.2)=2.16

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.16*1.05)/(0.15-0.05)

=22.68

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.7/1.15+2.16/1.15^2+22.68/1.15^2

=$21.13(Approx).